OpenStudy (anonymous):

Row Reduce {{-7,4,2,0},{4,-7,-2,0},{0,0,0,0}} I know the answer is {{1,0,-2/11,0},{0,1,2/11,0},{0,0,0,0}} but please show me the steps to get there.

OpenStudy (anonymous):

\[\left( \begin{array}{cccc} -7 & 4 & 2 & 0 \\ 4 & -7 & -2 & 0 \\ 0 & 0 & 0 & 0 \end{array} \right)\] this one?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[ \left[\begin{matrix} -7 & 4 & 2 & 0 \\ 4 & -7 & -2 & 0 \\ 0 & 0 & 0 & 0 \end{matrix}\right] \cong \left[\begin{matrix} -7 & 4 & 2 & 0 \\ -3 & -3 & 0& 0 \\ 0 & 0 & 0 & 0 \end{matrix}\right]\cong \left[\begin{matrix} -7 & 4 & 2 & 0 \\ 1 & 1 & 0& 0 \\ 0 & 0 & 0 & 0 \end{matrix}\right]\]\[\cong \left[\begin{matrix} 1 & 1 & 0& 0 \\ -7 & 4 & 2& 0 \\ 0 & 0 & 0 & 0 \end{matrix}\right]\cong \left[\begin{matrix} 1 & 1 & 0& 0 \\ 0 & 11 & 2& 0 \\ 0 & 0 & 0 & 0 \end{matrix}\right]\]

OpenStudy (anonymous):

Operationally: R1+R2 -> R2 (-1/3)R2 -> R2 R1 <-> R2 7R1 + R2 -> R2

OpenStudy (anonymous):

The trick is to look for things you can make 0.

OpenStudy (anonymous):

Funny, those are the exact steps I did, the only problem is I am having trouble getting it to row reduced echelon form. I'll figure it out, eventually. Thanks for your help.

OpenStudy (anonymous):

Oh, you need reduced row

OpenStudy (anonymous):

I got it! Thanks

OpenStudy (anonymous):

\[\cong\left[\begin{matrix} 1 & 0 & -\frac{2}{11} & 0 \\ 0 & 1 & \frac{2}{11}& 0 \\ 0 & 0 & 0 & 0 \end{matrix}\right]\]

OpenStudy (anonymous):

yup! I solved the problem already. Thanks for you help

OpenStudy (anonymous):

Great job.

OpenStudy (anonymous):

You too.